Determine the equation of g(x) that results from translating the function f(x) = (x + 10)^2 to the right 15 units.
g(x) = (x - 5)^2
g(x) = (x + 25)^2
g(x) = (x + 10)^2 – 15
g(x) = (x + 10)^2 + 15
I say g(x)=(x+10)^2-15 is the correct answer. Is this correct?

Answer by Alan3354(69443)

Determine the equation of g(x) that results from translating the function f(x) = (x + 10)^2 to the right 15 units.
g(x) = (x - 5)^2
g(x) = (x + 25)^2
g(x) = (x + 10)^2 – 15
g(x) = (x + 10)^2 + 15
I say g(x)=(x+10)^2-15 is the correct answer. Is this correct?
g(x)=(x+10)^2-15 moves it down 15
------
It's g(x) = (x+10 - 15)^2
= (x - 5)^2